OpenStudy (anonymous):

Write the sum as a product. Simplify the product. (–2) + (–2) + (–2) + (–2

OpenStudy (anonymous):

\[(-2)+1\times(-2)+1\times(-2)+1\times(-2)\]\[=-2-2-2-2\]\[=?\]solve it @mrs.dejoie1